# A geometric approximation of non-local interface and boundary conditions

### v1

@@ Abstract.
$\delta^{\prime}$-type interactions; non-local interface conditions; homogenization; Riemmanian manifold; Neumann sieve; resolvent convergence; spectrum; varying Hilbert spaces
@@ 1. Introduction
While these models are both useful and mathematically tractable, it is important to remember that the singular interaction represents an idealized approximation of a more realistic physical description. Consequently, a central problem in this area is to understand how such interactions can be approximated by regular models – for example, by Schrödinger operators with smooth potentials supported in shrinking neighborhoods of the surface $\Gamma$. Such approximations are well known for $\delta$-interactions (see [10, 19, 20, 6, 38]), whereas for $\delta^{\prime}$-interactions, a result on approximation by regular Schrödinger operators is known only in the one-dimensional case in which $\Gamma$ is a point and the normal derivatives in (1.1) are replaced by the usual derivative – see [4, 21] for the result with full mathematical rigor following a seminal physicists’ idea [16]. However, also another kind of regular approximations of $\delta^{\prime}$-interactions is known – as a homogenization limit of the Neumann Laplacian on a suitable perforated domain $\Omega_{\varepsilon}$. This domain has the form $\Omega\setminus\Gamma_{\varepsilon}$, where $\Omega$ is a domain on which our operator of interest is given (it is intersected by $\Gamma$), while $\Gamma_{\varepsilon}$ is either a subset of $\Gamma$ obtained by drilling a lot of small holes in it (thin sieve) or an $\varepsilon$-neighbourhood of $\Gamma$ punctured by many narrow passages (thick sieve [^1] The term ’thick sieve’ may seem misleading, as the sieve thickness tends to zero as $\varepsilon\to 0$. It was introduced in [18], presumably to distinguish this setting from the earlier studies of sieves with zero thickness.), see Figure 1.1. As $\varepsilon\to 0$, the number of holes (respectively, passages) tends to infinity, while their diameters tend to zero. It is known that, if the holes (respectively, passages) are appropriately scaled, the limiting behavior yields the desired $\delta^{\prime}$-interaction – see [8, 17, 31, 33, 34, 36] (respectively, [18, 32]). Needless to say, such geometric approximations are only possible when the approximated operator is non-negative (since any Neumann Laplacian is by definition a non-negative operator); this is the case when the interaction strength in (1.1) is a non-negative function.
Figure 1.1. The domain $\Omega$ perforated by a thin sieve (left image) and a thick sieve (right image). Under suitable assumptions on the scalings of the sieves, the Neumann Laplacian on this domain approximates the Laplacian on $\Omega$ subject to the ’standard’ $\delta^{\prime}$-type conditions (1.1).
@@
with a non-negative symmetric continuous kernel $K(x,y)$; we denote this operator by $\mathscr{H}$. Note that if we (formally) put $K(x,y)=\gamma(x)\delta(x-y)$ in (1.2), where $\delta(x-y)$ is a Dirac $\delta$-function on $\Gamma\times\Gamma$ supported on $\{(x,y)\in\Gamma\times\Gamma:x=y\}$, we arrive at the ’standard’ $\delta^{\prime}$-interaction (1.1). The goal of this paper is to construct approximations of this operator by Neumann Laplacians.
The nonlocal nature of the interface conditions (1.2) suggests to consider a thick sieve, but with passages connecting general points, not just mirror images on the two sieve borders. This concept has a natural physical appeal. If two regions are connected through a thick ’porous’ interface, the passages are rarely straight unless the sieve material is a product of a sophisticated engineering. Implementing this ideas mathematically, however, we immediately face difficulties. First of all, in the two-dimensional case, intersections between the passages become then unavoidable, whereas in dimensions $n\geq 3$, although they can be avoided, we obtain in general a highly complex sieve geometry. Secondly, if we allow the passages to connect distant points, their lengths do not vanish in the limit – even though they lie within a shrinking neighborhood of $\Gamma$. However, to get $\delta^{\prime}$-type interactions the passages must be short (cf. [18, 32]).
@@
Here $\Gamma$ is a subset of $\partial\Omega$, $\nu$ is the outward unit normal to $\partial\Omega$. The analysis of Laplacians with abstract non-local boundary conditions (covering those mentioned above) has been carried out in [25, 13]. Our objective remains the same: to approximate such operators by Neumann Laplacians. In this case, the approximating manifold consists of a domain $\Omega$ connected by numerous short thin ’handles’ – see Theorem 5.1 for precise statement of the result.
Elliptic operators with non-local boundary conditions frequently serve as infinitesimal generators of Feller semigroups associated with diffusion processes. These non-local formulations extend classical models by capturing boundary behaviors such as jumps, surface diffusion, or interactions beyond the local framework – see the pioneering works [22, 23, 40] and some key follow-ups [37, 24, 39]. In particular, diffusion equations with non-local Robin conditions resembling those in (1.3) were addressed in [7]. In this context, our homogenization result (cf. (5.2) in Theorem 5.1) provides an interpretation of the boundary conditions (1.3): a diffusing particle reaches the boundary from which it returns to the bulk, but due to the presence of attached passages connecting distant points, its motion continues starting from a different location.
The homogenization problem considered in this paper is closely related to the homogenization of Laplacians on Riemannian manifolds consisting of a base manifold with numerous attached ’handles’ or (in the case of a multicomponent base manifold) ’wormholes’ connecting the base manifold components. Such problems have been studied in [29, 30, 14, 5]. The key difference between the manifolds analyzed in those works and the one we consider here lies in the placement of the handles and wormholes: in previous studies, they are attached throughout the interior of the manifold, whereas in our model, they are attached along the boundary. In both settings, the homogenized operator may be non-local (cf. [29, 14]). However, in contrast to our case – where non-locality arises in the boundary conditions – the non-local terms in the cited works appear in the operator’s action itself.
The paper is organized as follows. In Section 2, we rigorously define the Laplacian with the interface conditions (1.2), construct the approximating manifold, and formulate the main result (Theorem 2.3) which establishes the resolvent convergence of the associated operators. As a consequence, we also state the convergence of eigenvalues and eigenfunctions (Theorem 2.5) and the convergence of the semigroups (Theorem 2.6). Section 3 presents an example of a manifold that satisfies the assumptions we claim and yields an arbitrary prescribed integral kernel $K(x,y)$ in the homogenization limit. The proofs of Theorems 2.32.6 are provided in Section 4. Finally, in Section 5, we discuss the approximation of the Laplacian with Robin-type non-local boundary conditions.
@@ 2. Problem setting and main results
Our main goal is to approximate the resolvent $(\mathscr{H}+\mathrm{Id})^{-1}$ of the operator $\mathscr{H}$ by the resolvent $(\mathscr{H}_{\varepsilon}+\mathrm{Id})^{-1}$ of the Neumann Laplacian $\mathscr{H}_{\varepsilon}$ on an appropriately constructed ’sieve-type’ Riemannian manifold $M_{\varepsilon}$; hereinafter, by $\mathrm{Id}$ we denote the identity operator. Note that $u=(\mathscr{H}+\mathrm{Id})^{-1}f$ with $f\in L^{2}(\Omega)$ is the weak solution to the following boundary value problem:
$\begin{cases}-\Delta u^{\pm}+u^{\pm}=f^{\pm}&\text{in }\Omega^{\pm},\\ \displaystyle\frac{\partial u^{\pm}}{\partial x^{n}}\mp\int_{\Gamma}K(x,y)(u^{\pm}(x)-u^{\mp}(y)){\,\mathrm{d}s_{y}}=0&\text{on }\Gamma,\\[5.69054pt] \displaystyle\frac{\partial u}{\partial\nu}=0&\text{on }\partial\Omega.\end{cases}$
@@
$T_{i,j,\varepsilon}$$\coloneqq\{x\in\mathbb{R}^{n}:(x^{1},\dots,x^{n-1})\in\mathbf{D}_{i,\varepsilon},\,|x^{n}|\leq h_{i,j,\varepsilon}/2\}$
$\,\cong\{x\in\mathbb{R}^{n}:(x^{1},\dots,x^{n-1})\in\mathbf{D}_{j,\varepsilon},\,|x^{n}|\leq h_{i,j,\varepsilon}/2\},$ (2.6)
@@
$S_{i,\varepsilon}^{+}$$\coloneqq\{x\in\mathbb{R}^{n}:(x^{1},\dots,x^{n-1})\in\mathbf{D}_{i,\varepsilon},\,x^{n}=h_{i,j,\varepsilon}/2\},$
$S_{j,\varepsilon}^{-}$$\coloneqq\{x\in\mathbb{R}^{n}:(x^{1},\dots,x^{n-1})\in\mathbf{D}_{i,\varepsilon},\,x^{n}=-h_{i,j,\varepsilon}/2\},\ j=\ell_{\varepsilon}(i).$
Finally, for $(i,j)\in\mathbb{L}_{\varepsilon}$, we glue the set $T_{i,j,\varepsilon}$ to $M^{+}$ (respectively, $M^{-}$) by identifying $S^{+}_{i,\varepsilon}$ and $D^{+}_{i,\varepsilon}$ (respectively, $S^{-}_{i,\varepsilon}$ and $D^{-}_{j,\varepsilon}$). By construction, the sets $S^{\pm}_{i,\varepsilon}$ and $D^{\pm}_{j,\varepsilon}$ are congruent, and the above identification is carrying out according to this congruence. In this way, we obtain the (topological) manifold
@@
$\widehat{\mathbf{D}}^{\pm}_{i,\varepsilon}\coloneqq(d_{i,\varepsilon}^{\pm})^{-1}\mathbf{D}_{i,\varepsilon}^{\pm}$
chosen in such a way that the outerradius of $\widehat{\mathbf{D}}_{i,\varepsilon}$ equals $1$. In the following, the symbol $\Lambda_{\rm N}(\Omega)$ means the smallest non-zero eigenvalue of the Neumann Laplacian on a bounded connected open set $\Omega\subset\mathbb{R}^{d}$ with Lipschitz boundary. Our assumptions read as follows,
@@
$\mathscr{B}_{n-1}({\varrho^{\pm}_{i,\varepsilon}},\mathbf{x}^{\pm}_{i,\varepsilon})\cap\mathscr{B}_{n-1}(\varrho^{\pm}_{j,\varepsilon},\mathbf{x}^{\pm}_{j,\varepsilon})=\emptyset,\;\forall i,j\in\mathbb{I}_{\varepsilon},\ i\not=j,$ (2.11)
@@
To formulate the last two assumptions, we introduce the quantities $K_{i,j,\varepsilon}$, $(i,j)\in\mathbb{L}_{\varepsilon}$ by
$K_{i,j,\varepsilon}\coloneqq\operatorname{vol}_{n-1}(\mathbf{D}_{i,\varepsilon})h_{i,j,\varepsilon}^{-1}$ (2.16)
@@
$K_{i,j,\varepsilon}h_{i,j,\varepsilon}(\varrho_{i,\varepsilon}^{\pm})^{1-n}=\operatorname{vol}_{n-1}(\mathbf{D}_{i,\varepsilon}^{\pm})(\varrho_{i,\varepsilon}^{\pm})^{1-n}=\operatorname{vol}_{n-1}(\widehat{\mathbf{D}}_{i,\varepsilon}^{\pm})(d_{i,\varepsilon}^{\pm})^{n-1}(\varrho_{i,\varepsilon}^{\pm})^{1-n}\geq C(d_{i,\varepsilon}^{\pm})^{n-1}(\varrho_{i,\varepsilon}^{\pm})^{1-n},$
where at the last step we use (2.8). Hence, due to (2.15) and (2.17), we get
$\sup_{i\in\mathbb{I}_{\varepsilon}}d_{i,\varepsilon}^{\pm}(\varrho_{i,\varepsilon}^{\pm})^{-1}\leq Ch_{\varepsilon}\to 0\,\text{ as }\,\varepsilon\to 0.$ (2.19)
@@
The dependence of this operator on $\varepsilon$ is merely formal – it just ’drags’ $f$ from $M^{+}\cup M^{-}$ to $\Omega^{+}\cup\Omega^{-}$, while the values of $f$ on $T_{i,j,\varepsilon}$ do not affect $\mathscr{J}_{\varepsilon}f$. Note that $\mathscr{J}_{\varepsilon}f\in H^{1}(\Omega\setminus\Gamma)$ provided $f\in H^{1}(M_{\varepsilon})$.
@@
Hence, Theorem 2.3 establishes a kind of strong resolvent convergence of the operators $\mathscr{H}_{\varepsilon}$ to operator $\mathscr{H}$ as $\varepsilon$ to 00.

### v2

@@ Abstract.
$\delta^{\prime}$-type interactions; non-local interface conditions; homogenization; Riemannian manifold; Neumann sieve; resolvent convergence; spectrum; varying Hilbert spaces
@@ 1. Introduction
While these models are both useful and mathematically tractable, it is important to remember that the singular interaction represents an idealized approximation of a more realistic physical description. Consequently, a central problem in this area is to understand how such interactions can be approximated by regular models – for example, by Schrödinger operators with smooth potentials supported in shrinking neighborhoods of the surface $\Gamma$. Such approximations are well known for $\delta$-interactions (see [10, 19, 20, 6, 38]), whereas for $\delta^{\prime}$-interactions, a result on approximation by regular Schrödinger operators is known only in the one-dimensional case in which $\Gamma$ is a point and the normal derivatives in (1.1) are replaced by the usual derivative – see [4, 21] for the result with full mathematical rigor following a seminal physicists’ idea [16]. However, also another kind of regular approximations of $\delta^{\prime}$-interactions is known – as a homogenization limit of the Neumann Laplacian on a suitable perforated domain $\Omega_{\varepsilon}$. This domain has the form $\Omega\setminus\Gamma_{\varepsilon}$, where $\Omega$ is a domain on which our operator of interest is given (it is intersected by $\Gamma$), while $\Gamma_{\varepsilon}$ is either a subset of $\Gamma$ obtained by drilling a lot of small holes in it (thin sieve) or an $\varepsilon$-neighbourhood of $\Gamma$ punctured by many narrow passages (thick sieve [^1] The term ‘thick sieve’ may seem misleading, as the sieve thickness tends to zero as $\varepsilon\to 0$. It was introduced in [18], presumably to distinguish this setting from the earlier studies of sieves with zero thickness.), see Figure 1.1. As $\varepsilon\to 0$, the number of holes (respectively, passages) tends to infinity, while their diameters tend to zero. It is known that, if the holes (respectively, passages) are appropriately scaled, the limiting behavior yields the desired $\delta^{\prime}$-interaction – see [8, 17, 31, 33, 34, 36] (respectively, [18, 32]). Needless to say, such geometric approximations are only possible when the approximated operator is non-negative (since any Neumann Laplacian is by definition a non-negative operator); this is the case when the interaction strength in (1.1) is a non-negative function.
Figure 1.1. The domain $\Omega$ perforated by a thin sieve (left image) and a thick sieve (right image). Under suitable assumptions on the scalings of the sieves, the Neumann Laplacian on this domain approximates the Laplacian on $\Omega$ subject to the ‘standard’ $\delta^{\prime}$-type conditions (1.1).
@@
with a non-negative symmetric continuous kernel $K(x,y)$; we denote this operator by $\mathscr{H}$. Note that if we (formally) put $K(x,y)=\gamma(x)\delta_{x}(y)$ in (1.2), where $\delta_{x}(y)$ is a Dirac $\delta$-function on $\Gamma$ supported at $x\in\Gamma$, we arrive at the ‘standard’ $\delta^{\prime}$-interaction (1.1). The goal of this paper is to construct approximations of this operator by Neumann Laplacians.
The nonlocal nature of the interface conditions (1.2) suggests to consider a thick sieve, but with passages connecting general points, not just mirror images on the two sieve borders. This concept has a natural physical appeal. If two regions are connected through a thick ‘porous’ interface, the passages are rarely straight unless the sieve material is a product of a sophisticated engineering. Implementing this ideas mathematically, however, we immediately face difficulties. First of all, in the two-dimensional case, intersections between the passages become then unavoidable, whereas in dimensions $n\geq 3$, although they can be avoided, we obtain in general a highly complex sieve geometry. Secondly, if we allow the passages to connect distant points, their lengths do not vanish in the limit – even though they lie within a shrinking neighborhood of $\Gamma$. However, to get $\delta^{\prime}$-type interactions the passages must be short (cf. [18, 32]).
@@
Here $\Gamma$ is a subset of $\partial\Omega$, $\nu$ is the outward unit normal to $\partial\Omega$. The analysis of Laplacians with abstract non-local boundary conditions (covering those mentioned above) has been carried out in [25, 13]. Our objective remains the same: to approximate such operators by Neumann Laplacians. In this case, the approximating manifold consists of a domain $\Omega$ connected by numerous short thin ‘handles’ – see Theorem 5.1 for precise statement of the result.
Elliptic operators with non-local boundary conditions frequently serve as infinitesimal generators of Feller semigroups associated with diffusion processes – see the pioneering works [22, 23, 40] and some key follow-ups [37, 24, 39]. In particular, diffusion equations with non-local Robin conditions resembling those in (1.3) were addressed in [7]. In this context, our homogenization result (cf. (5.2) in Theorem 5.1) provides an interpretation of the boundary conditions (1.3): a diffusing particle reaches the boundary from which it returns to the bulk, but due to the presence of attached passages connecting distant points, its motion continues starting from a different location.
The homogenization problem considered in this paper is closely related to the homogenization of Laplacians on Riemannian manifolds consisting of a base manifold with numerous attached ‘handles’ or (in the case of a multicomponent base manifold) ‘wormholes’ connecting the base manifold components. Such problems have been studied in [29, 30, 14, 5]. The key difference between the manifolds analyzed in those works and the one we consider here lies in the placement of the handles and wormholes: in previous studies, they are attached throughout the interior of the manifold, whereas in our model, they are attached along the boundary. In both settings, the homogenized operator may be non-local (cf. [29, 14]). However, in contrast to our case – where non-locality arises in the boundary conditions – the non-local terms in the cited works appear in the operator’s action itself.
The paper is organized as follows. In Section 2, we rigorously define the Laplacian with the interface conditions (1.2), construct the approximating manifold, and formulate the main result (Theorem 2.3) which establishes the resolvent convergence of the associated operators. As a consequence, we also state the convergence of eigenvalues and eigenfunctions (Theorem 2.5) and the convergence of the semigroups (Theorem 2.6). Section 3 presents an example of a manifold that satisfies the assumptions we claim and yields an arbitrary prescribed integral kernel $K(x,y)$ in the homogenization limit. The proofs of Theorems 2.3, 2.5, 2.6 are provided in Section 4. Finally, in Section 5, we discuss the approximation of the Laplacian with Robin-type non-local boundary conditions.
@@ 2. Problem setting and main results
Our main goal is to approximate the resolvent $(\mathscr{H}+\mathrm{Id})^{-1}$ of the operator $\mathscr{H}$ by the resolvent $(\mathscr{H}_{\varepsilon}+\mathrm{Id})^{-1}$ of the Neumann Laplacian $\mathscr{H}_{\varepsilon}$ on an appropriately constructed ‘sieve-type’ Riemannian manifold $M_{\varepsilon}$; hereinafter, by $\mathrm{Id}$ we denote the identity operator. Note that $u=(\mathscr{H}+\mathrm{Id})^{-1}f$ with $f\in L^{2}(\Omega)$ is the weak solution to the following boundary value problem:
$\begin{cases}-\Delta u^{\pm}+u^{\pm}=f^{\pm}&\text{in }\Omega^{\pm},\\ \displaystyle\frac{\partial u^{\pm}}{\partial x^{n}}=\pm\int_{\Gamma}K(x,y)(u^{\pm}(x)-u^{\mp}(y)){\,\mathrm{d}s_{y}}&\text{on }\Gamma,\\[5.69054pt] \displaystyle\frac{\partial u}{\partial\nu}=0&\text{on }\partial\Omega.\end{cases}$
@@
$T_{i,j,\varepsilon}$$\coloneqq\{x\in\mathbb{R}^{n}:(x^{1},\dots,x^{n-1})\in\mathbf{D}_{i,\varepsilon}^{+},\,|x^{n}|\leq h_{i,j,\varepsilon}/2\}$
$\,\cong\{x\in\mathbb{R}^{n}:(x^{1},\dots,x^{n-1})\in\mathbf{D}_{j,\varepsilon}^{-},\,|x^{n}|\leq h_{i,j,\varepsilon}/2\},$ (2.6)
@@
$S_{i,\varepsilon}^{+}$$\coloneqq\{x\in\mathbb{R}^{n}:(x^{1},\dots,x^{n-1})\in\mathbf{D}^{+}_{i,\varepsilon},\,x^{n}=h_{i,j,\varepsilon}/2\},$
$S_{j,\varepsilon}^{-}$$\coloneqq\{x\in\mathbb{R}^{n}:(x^{1},\dots,x^{n-1})\in\mathbf{D}^{+}_{i,\varepsilon},\,x^{n}=-h_{i,j,\varepsilon}/2\},\ j=\ell_{\varepsilon}(i).$
Finally, for $(i,j)\in\mathbb{L}_{\varepsilon}$, we glue the set $T_{i,j,\varepsilon}$ to $M^{+}$ (respectively, $M^{-}$) by identifying $S^{+}_{i,\varepsilon}$ and $D^{+}_{i,\varepsilon}$ (respectively, $S^{-}_{i,\varepsilon}$ and $D^{-}_{j,\varepsilon}$). By construction, the sets $S^{+}_{i,\varepsilon}$, $D^{+}_{i,\varepsilon}$, $S^{-}_{j,\varepsilon}$, $D^{-}_{j,\varepsilon}$ ($(i,j)\in\mathbb{L}_{\varepsilon}$) are congruent, and the above identification is carrying out according to this congruence. In this way, we obtain the (topological) manifold
@@
$\widehat{\mathbf{D}}^{\pm}_{i,\varepsilon}\coloneqq(d_{i,\varepsilon}^{\pm})^{-1}\mathbf{D}_{i,\varepsilon}^{\pm}.$
Evidently, the outerradius of $\widehat{\mathbf{D}}_{i,\varepsilon}$ equals $1$. In the following, the symbol $\Lambda_{\rm N}(\Omega)$ means the smallest non-zero eigenvalue of the Neumann Laplacian on a bounded connected open set $\Omega\subset\mathbb{R}^{d}$ with Lipschitz boundary. Our assumptions read as follows,
@@
$\mathscr{B}_{n-1}({\varrho^{\pm}_{i,\varepsilon}},\mathbf{x}^{\pm}_{i,\varepsilon})\cap\mathscr{B}_{n-1}(\varrho^{\pm}_{k,\varepsilon},\mathbf{x}^{\pm}_{k,\varepsilon})=\emptyset,\;\forall i,k\in\mathbb{I}_{\varepsilon},\ i\not=k,$ (2.11)
@@
To formulate the last two assumptions, for $(i,j)\in\mathbb{L}_{\varepsilon}$ we introduce the quantities $K_{i,j,\varepsilon}$ by
$K_{i,j,\varepsilon}\coloneqq\operatorname{vol}_{n-1}(\mathbf{D}_{i,\varepsilon}^{+})h_{i,j,\varepsilon}^{-1}=\operatorname{vol}_{n-1}(\mathbf{D}_{j,\varepsilon}^{-})h_{i,j,\varepsilon}^{-1},\quad(i,j)\in\mathbb{L}_{\varepsilon}$ (2.16)
@@
$\forall i\in\mathbb{I}_{\varepsilon}:\quad h_{i,j,\varepsilon}K_{i,j,\varepsilon}(\varrho_{i,\varepsilon}^{+})^{1-n}$ $=\operatorname{vol}_{n-1}(\mathbf{D}_{i,\varepsilon}^{+})(\varrho_{i,\varepsilon}^{+})^{1-n}=\operatorname{vol}_{n-1}(\widehat{\mathbf{D}}_{i,\varepsilon}^{+})(d_{i,\varepsilon}^{+})^{n-1}(\varrho_{i,\varepsilon}^{+})^{1-n}$
$\geq C(d_{i,\varepsilon}^{+})^{n-1}(\varrho_{i,\varepsilon}^{+})^{1-n},\text{ where }j=\ell_{\varepsilon}(j).$ (2.19)
where at the last step we use (2.8). Similarly, one has
$\forall j\in\mathbb{I}_{\varepsilon}:\quad h_{i,j,\varepsilon}K_{i,j,\varepsilon}(\varrho_{j,\varepsilon}^{-})^{1-n}\geq C(d_{j,\varepsilon}^{-})^{n-1}(\varrho_{j,\varepsilon}^{-})^{1-n},\text{ where }i=\ell_{\varepsilon}^{-1}(j).$ (2.20)
From (2.15), (2.17), (2.19), (2.20) we deduce
$\sup_{i\in\mathbb{I}_{\varepsilon}}d_{i,\varepsilon}^{\pm}(\varrho_{i,\varepsilon}^{\pm})^{-1}\leq\sqrt[n-1]{h_{\varepsilon}}\,\to 0\,\text{ as }\,\varepsilon\to 0.$ (2.21)
@@
The dependence of this operator on $\varepsilon$ is merely formal – it just ‘drags’ $f$ from $M^{+}\cup M^{-}$ to $\Omega^{+}\cup\Omega^{-}$, while the values of $f$ on $T_{i,j,\varepsilon}$ do not affect $\mathscr{J}_{\varepsilon}f$. Note that $\mathscr{J}_{\varepsilon}f\in H^{1}(\Omega\setminus\Gamma)$ provided $f\in H^{1}(M_{\varepsilon})$.
@@
Hence, Theorem 2.3 establishes a kind of strong resolvent convergence of the operators $\mathscr{H}_{\varepsilon}$ to operator $\mathscr{H}$ as $\varepsilon\to 0$.